# Direct Alignment of Draft Model for Speculative Decoding with Chat-Fine-tuned LLMs

### v3

@@ 2.1 Draft model pretraining
A draft model is required to have the same tokenizer as the corresponding target model with the same or a subset of vocabulary. Some language model families sharing the same tokenizer have wide ranges of model sizes where we can pick draft models at a desired sizes, e.g., Pythia (biderman2023pythia). However, this is not the case more often, e.g., Llama 2 (Touvron et al., 2023), where the smallest is still too large. In such cases, we train a draft language model from scratch with the regular next token prediction loss in order to have a good language model to be aligned with the target model more effectively. It is possible to incorporate initialization techniques such as Weight Subcloning (samragh2023weight) to expedite this step. While it may be trivial, in our empirical observations, pretrained draft models show significantly better alignment to target model, compared to those without pretraining.
@@ 3 Experiments
Evaluation metrics and tasks. We evaluate our draft model by measuring the (1) block efficiency ($\tau$): average number of tokens generated per block (or target model run), for a block of size $\gamma$ and input $x$, the maximum value of $\tau(x)$ can be $\gamma+1$ and (2) memory-bound speed-up (MBSU): hypothetical speed-up achieved by SD for a given block efficiency $\tau(x)$ and a relative latency $c$ defined as ratio between number of parameters of draft to target model, i.e., $\mathrm{MBSU}(x):=\frac{c\tau(x)}{c\gamma+1}$. We measure these metrics on a variety of tasks using different block lengths $\gamma$ in $\{3,5\}$. The evaluation tasks include: (a) open-ended text generation (Conover et al., 2023, databricks-Dolly-15k), (b) extreme summarization (Narayan et al., 2018, XSum), and (c) news summarization (Nallapati et al., 2016, CNN-dailymail). For open-ended text generation we random-sample with temperature$=0.6$ and top-$p$$=0.9$ for both draft model and target model, while we greedy-sample for the rest.
@@
Figure 1: Draft models evaluated on Memory-Bound Speed-Up (MBSU) for multiple tasks (Dolly, CNN-DM, XSum) with draft lengths (3, 5) and training losses (KLD, TVD, TVD++);

### v4

@@ 2.1 Draft model pretraining
A draft model is required to have the same tokenizer as the corresponding target model with the same or a subset of vocabulary. Some language model families sharing the same tokenizer have wide ranges of model sizes where we can pick draft models at a desired sizes, e.g., Pythia (Biderman et al., 2023). However, this is not the case more often, e.g., Llama 2 (Touvron et al., 2023), where the smallest is still too large. In such cases, we train a draft language model from scratch with the regular next token prediction loss in order to have a good language model to be aligned with the target model more effectively. It is possible to incorporate initialization techniques such as Weight Subcloning (Samragh et al., 2023) to expedite this step. While it may be trivial, in our empirical observations, pretrained draft models show significantly better alignment to target model, compared to those without pretraining.
@@ 3 Experiments
Evaluation metrics and tasks. We evaluate our draft model by measuring the (1) block efficiency ($\tau$): average number of tokens generated per block (or target model run), for a block of size $\gamma$ and input $x$, the maximum value of $\tau(x)$ can be $\gamma+1$, (2) memory-bound speed-up (MBSU): hypothetical speed-up achieved by SD for a given block efficiency $\tau(x)$ and a relative latency $c$ defined as ratio between number of parameters of draft to target model, i.e., $\mathrm{MBSU}(x):=\frac{c\tau(x)}{c\gamma+1}$, and (3)Token Rate: number of tokens generated per second, in our case we measure the ratio of the SPD token-rate to auto-regressive token-rate, a ratio $>1$ implies higher number of tokens were generated per second for SPD. We measure these metrics on a variety of tasks using different block lengths $\gamma$ in $\{3,5\}$. The evaluation tasks include: (a) open-ended text generation (Conover et al., 2023, databricks-Dolly-15k), (b) extreme summarization (Narayan et al., 2018, XSum), and (c) news summarization (Nallapati et al., 2016, CNN-dailymail). For open-ended text generation we random-sample with temperature$=0.6$ and top-$p$$=0.9$ for both draft model and target model, while we greedy-sample for the rest.
@@
Figure 1: Draft models evaluated on Memory-Bound Speed-Up (MBSU) and token-rate (realtive to auto-regressive generation) for multiple tasks (Dolly, CNN-dailymail, XSum) with draft lengths (3, 5) and training losses (KLD, TVD, TVD++);